OpenStudy (anonymous):

For \(n \geq 1\) set \( f(x) = \frac{1}{(1-x)^{n+1} }\) . Calculate the power series expansion of \(f\) at \(x_{0} = 0\) by two different ways: a) Apply the formula the binomial series.

OpenStudy (anonymous):

for any n?

OpenStudy (anonymous):

i think yes..

OpenStudy (anonymous):

ok maybe it is not that bad

OpenStudy (anonymous):

we have \[\frac{1}{1-x}=\sum_{k=0}^{\infty}x^k\] then \[\frac{1}{(1-x)^2}\] is the derivative of the previous one so it is \[\sum_{k=1}^\infty kx^{k-1}\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

if you want to keep starting at \(k=0\) you can write it as \[\sum_{k=0}^{\infty}(k+1)x^k\] \[\frac{1}{(1-x)^3}\] is almost the derivative of the previous one, but not quite , it is \(\frac{1}{2}\) times that one so you get \[\frac{1}{(1-x)^3}=\sum_{k=0}^{\infty}\frac{1}{2}(n+1)(n+2)x^n\]

OpenStudy (anonymous):

i don't know the k became an n

OpenStudy (anonymous):

but you get the idea i guess you can arrive at a general formula but i am too befuddled to do it can probably google it if you like

OpenStudy (anonymous):

ok thanks for your effort satellite i made you very tired today :)

OpenStudy (anonymous):

yeah a quick search gave me this http://www.ucl.ac.uk/Mathematics/geomath/level2/series/ser8.html you might come up with a better one, but i think the idea is right. take successive derivatives and adjust

OpenStudy (anonymous):

oh ignore that it is not what you are doing sorry

OpenStudy (anonymous):

i mean ignore the link

OpenStudy (anonymous):

ok no probleem

OpenStudy (anonymous):

thank you satellite